I need your help with solving equations as relations. The problem is as follows:
3x=2y - 1 a. (1,2) b.(-1,-1) c.(-2,-5/2) d.(0,-1/2)
I tried and this is what i got:
3x=2y - 1 Add one to the 3x
3x + 1=2y You get 4x = 2y
4x = 2y Divide 4x by 2y and you get 2xy.
I dont think I did that right and thats not one of my multiple choice answers. Thanks,
Addy

Answer by stanbon(48558)

To find the correct answer you need to substitute
each of the (x,y) pairs into the given equation to
see which pair gives you a true statement.
Trying (1,2) you get 3(1)=2(2)-1 or 3=3.
(1,2) gives you a true statement.
You then try (-1,-1); it also works.
So does (-2,-5.2)
Only (0,-1/2) does not give a true statement.
Maybe you were asked to find the pair that was NOT in the relation.
Cheers,
Stan H.